OpenStudy (anonymous):

4. Which of the following solutions has the highest number of sodium ions? The answer is 50.00 mL of 0.125M NaCl but I'm not sure how to calculate it

OpenStudy (somy):

what are other options?

OpenStudy (somy):

i'll show you how to calculate it using other options and then you will try with the one you are saying is the answer

OpenStudy (anonymous):

Here's all the options 20.00 mL of 0.200M NaCl 25.00 mL of 0.175M NaCl 30.00 mL of 0.145M NaCl 50.00 mL of 0.125M NaCl 100.00 mL of 0.0500M NaCl

OpenStudy (somy):

which one do you want me to try aside from the right one?

OpenStudy (anonymous):

Any one you'd like.

OpenStudy (somy):

lets do 30.00 mL of 0.145M NaCl

OpenStudy (somy):

so i'll start going stepwise i'll explain how to do and give formulas you will calculate agreed?

OpenStudy (anonymous):

Okay

OpenStudy (somy):

so the q is ' highest number of sodium ions ' for us to get there we have several steps now first to start when you look at whats given 30.00 mL of 0.145M NaCl what do you see? volume and concentration, right? now question for you what can you find if you have these two?

OpenStudy (anonymous):

I'm not completely sure. Sorry.

OpenStudy (somy):

alright you can find mole here is the formula concentration= mole / volume concentration unit is mole/dm3 which is basically M (molar) volume unit in the formula has to be dm3 do it and then use this formula and find mole

OpenStudy (anonymous):

0.145 M Nacl/ 30.00=.005?

OpenStudy (somy):

look carefully at unit of volume

OpenStudy (somy):

unit has to be dm3 while yours is mL

OpenStudy (anonymous):

Should it be 30.00^3?

OpenStudy (anonymous):

Or 30 x 3?

OpenStudy (somy):

nope 1000 mL = 1 dm3 30 mL = ? dm3

OpenStudy (anonymous):

.33 dm3?

OpenStudy (somy):

okay look 1000mL = 1 dm3 30mL = X dm3 so its \[1000mL \times X = 30mL \times 1 dm3 \]

OpenStudy (somy):

if you make X as subject it will be = 30/1000

OpenStudy (anonymous):

Ooops. I did 1000/30 at first. 30/1000=.03 So it's .03 dm3?

OpenStudy (somy):

yep)

OpenStudy (somy):

now solve it the formula

OpenStudy (anonymous):

0.145/.03 dm3=4.83 dm3?

OpenStudy (anonymous):

Or should the dm3 drop off and just leave 0.145/.03=4.83?

OpenStudy (somy):

nope look the formula is \[concentration = \frac{ mole }{ volume }\]

OpenStudy (somy):

if you make mole as subject its gonna be \[mole= concentration \times volume \] isnt it?

OpenStudy (anonymous):

Sorry. My computer is having some issues today and keeps shutting down on me. Okay now I'm confused. You said it wasn't 0.145/ 30.00 which is moles/volume

OpenStudy (somy):

0.145 M is not mole its concentration! mole is what you are supposed to find

OpenStudy (anonymous):

Is that not mole? it says M beside it which means mole

OpenStudy (somy):

no M is molar which is concentration! mole's unit is gram/amu pretty much

OpenStudy (somy):

its always written as mole though mole, and its unit as mole

OpenStudy (anonymous):

Okay. Got it. So I'm dividing what by .03 exactly?

OpenStudy (somy):

you are not dividing look at the formula

OpenStudy (anonymous):

It's mole over volume

OpenStudy (somy):

;yes but you are trying to find mole! so we have to rearrange the formula originally the formula is concentration = mole/ volume we need mole so we have to make mole as a subject of this formula mole = concentration * volume so now if we make mole as subject- we have to multiply concentration by volume

OpenStudy (anonymous):

So would that make it 0.145x.03?

OpenStudy (somy):

yes

OpenStudy (anonymous):

0.145x.03=.004

OpenStudy (somy):

yep

OpenStudy (somy):

so now u found mole of NaCl from here since we now have mole of the whole compound we can find mole of sodium alone!

OpenStudy (somy):

mole= mass/molecular mass 0.004 = total Mr X = Mr of sodium total mole * Mr of sodium / total Mr = mole of Sodium

OpenStudy (somy):

so im just using simple math if 0.004 mole is mole of whole NaCl what is mole of Na? since we know molecular masses of Na and Cl you can calculate the total molecular mass of NaCl compound and then take molecular mass of sodium alone so 0.004 mole = molecular mass of NaCl and X mole of sodium alone = molecular mass of sodium so if we do cross multiplication we will get 0.004mole * molecular mass of sodium / molecular mass of NaCl = X moles of sodium alone

OpenStudy (somy):

do you get it?

OpenStudy (anonymous):

So the molecular mass of sodium in 22.99 so it's 0.004x22.99/0.004?

OpenStudy (somy):

look carefully at the formula please

OpenStudy (somy):

0.004mole * molecular mass of sodium / molecular mass of NaCl = X moles of sodium alone its divided by molecular mass of NaCl

OpenStudy (anonymous):

You said that 0.004= molecular of NaCl. I got confused by that. The molecular mass of NaCl is 58.44 so the formula would be 0.004*22.99/58.44?

OpenStudy (somy):

that is mathematically we are making equation that way yes this one is correct

OpenStudy (anonymous):

The answer it 2.047

OpenStudy (somy):

good job now last step multiply mole of sodium by Avogadro's number and you will get number of sodium ions Avogadro's number is 6.022* 10^23

OpenStudy (anonymous):

2.047x6.22x10^23=12.732x10^23

OpenStudy (somy):

yep

OpenStudy (somy):

do same thing with the right answer and then compare these final answers see the difference for yourself

OpenStudy (anonymous):

0.125x.03=.00375 .00375x22.99/58.44=.0015 .0015x6.022x10^23=.009x10^23

OpenStudy (anonymous):

Okay what did I do wrong because that's smaller

OpenStudy (somy):

you are mixed up numbers

OpenStudy (somy):

volume in your right answer is 50mL! not 30mL like we did just now

OpenStudy (anonymous):

So the first step it 50/1000 which is ,05

OpenStudy (somy):

yep

OpenStudy (anonymous):

then 0.125x.05=.006

OpenStudy (anonymous):

.006*22.99/58.44=.0023

OpenStudy (anonymous):

*.0024

OpenStudy (anonymous):

When you multiply that by 6.022 x10^23 it comes out smaller too

OpenStudy (somy):

oh wait!

OpenStudy (somy):

0.004*22.99/58.44 you calculated it wrong!!!

OpenStudy (somy):

thats why it seems smaller

OpenStudy (somy):

0.0016 this is the answer

OpenStudy (somy):

multiply it by avogadro's number now it will be right

OpenStudy (somy):

sorry didnt notice your wrong answer as i was not calculating myself

OpenStudy (anonymous):

Oh hey. That'll do it

OpenStudy (somy):

all good now?

OpenStudy (anonymous):

So the answer to this one it .014

OpenStudy (anonymous):

Which is bigger. Finally.

OpenStudy (somy):

good)

OpenStudy (anonymous):

Fantastic. Thank you so much!

OpenStudy (somy):

you are welcome :))))